# Machine Learners Should Acknowledge the Legal Implications of Large Language Models as Personal Data

### v1

@@ Abstract
Does GPT know you? The answer depends on your level of public recognition; however, if your information was available on a website, the answer is probably yes. All Large Language Models (LLMs) memorize training data to some extent. If an LLM training corpus includes personal data, it also memorizes personal data. Developing an LLM typically involves processing personal data, which falls directly within the scope of data protection laws. If a person is identified or identifiable, the implications are far-reaching: the AI system is subject to EU General Data Protection Regulation requirements even after the training phase is concluded. To back our arguments: (1.) We reiterate that LLMs output training data at inference time, be it verbatim or in generalized form. (2.) We show that some LLMs can thus be considered personal data on their own. This triggers a cascade of data protection implications such as data subject rights, including rights to access, rectification, or erasure. These rights extend to the information embedded with-in the AI model. (3.) This paper argues that machine learning researchers must acknowledge the legal implications of LLMs as personal data throughout the full ML development lifecycle, from data collection and curation to model provision on, e.g., GitHub or Hugging Face. (4.) We propose different ways for the ML research community to deal with these legal implications. Our paper serves as a starting point for improving the alignment between data protection law and the technical capabilities of LLMs. Our findings underscore the need for more interaction between the legal domain and the ML community.
Machine Learning, ICML
@@ 1 Introduction
GPT probably knows you—not just you, but it has likely memorized data about countless other individuals as well (Verma et al., 2024). Most Large Language Models (LLMs) memorize parts of the input data provided during training in verbatim or quasi-verbatim manner (Carlini et al., 2022; Somepalli et al., 2023; Verma et al., 2024). This memorization is double-edged. The training objective of an LLM is partly generalization, but memorization is also an essential component (Tirumala et al., 2022; Power et al., 2022; Hartmann et al., 2023; Biderman et al., 2024). Without memorization, an LLM would be, for example, unable to tell us that the Eiffel Tower is located in Paris. Thus, memorization is an important and — to some extent — deliberate aspect of training LLMs. While memorization is a crucial feature of LLMs, it also comes with challenges (Bender et al., 2021), while a key problematic aspect is the consideration of memorized personal data.
@@
Although the problematic nature of personal data memorization in LLMs has been mentioned numerous times (Carlini et al., 2019; Hu et al., 2022; Carlini et al., 2022; Zhang et al., 2023a; Cooper & Grimmelmann, 2024), the debate about its implications is still in its early stages. While much of the technical work focuses on unlearning techniques (Juliussen et al., 2023; Bourtoule et al., 2021; Zhang et al., 2023b), the legal implications are often overlooked in the technical ML community [^1] We use the term “ML community” to refer to researchers — whether from academia or industry — who work on topics suitable for publication at conferences like ICML, NeurIPS, or ICLR.. For example, the processing of personal data requires a legal basis. Furthermore, individuals whose data has been memorized by an LLM may have the right to request its erasure from the model, which could require the removal of that information from the model itself, see discussion (Villaronga et al., 2018; Yaish, 2019; Fabbrini & Celeste, 2020; Allegri et al., 2022; Juliussen et al., 2023). These questions are already unfolding in real life. For instance, complaints have been filed against OpenAI, alleging General Data Protection Regulation (GDPR) [^2] EU Regulation 2016/679, 27.4.2016. violations due to the company’s failure to ensure the accuracy of personal data, comply with access and rectification requests, and provide transparency about the sources and processing of personal data in ChatGPT (NOBY, 4). From our perspective, not all model developers fully consider or acknowledge the serious implications of this requirement.
@@
Why should ML researchers care whether LLMs are personal data or not? If LLMs qualify as personal data, ML researchers may be responsible for their processing. In such cases, researchers must also comply with data protection regulations. As we demonstrate below, non-compliance with data protection regulations can lead to severe fines, whether on a personal, university, or company level. Please also note that some data protection laws, such as the GDPR, even extend to researchers outside the EU as long they process data from people within the EU (see Article 3 GDPR). Beyond legal compliance, successful LLM research (and its applications) requires not only technical rigor but also the trust of users. We believe that adhering to data protection regulations can enhance trust in ML systems. When researchers commit to data protection safeguards, their applications are more likely to gain public acceptance.
@@
Computational Perspective. LLMs clearly memorize some input data verbatim in normal usage contexts while still generalizing to new contexts (Carlini et al., 2019; Hartmann et al., 2023; Tirumala et al., 2022; Biderman et al., 2024). This memorization increases with the growing capacity of the model, the frequency of examples in the training data, and the number of tokens of context used to prompt the model (Carlini et al., 2022). If data are not outputted verbatim, in some settings, an adversarial attacker is able to extract large amounts of training data (Nasr et al., 2023a). However, it has also been noted that verbatim memorization of facts is an important feature — and not a bug — of LLMs (Biderman et al., 2024). A weaker form of verbatim memorization is the knowledge of facts in LLMs (Ippolito et al., 2022). LLMs store certain facts (Petroni et al., 2019; Hu et al., 2023; Roberts et al., 2020; Yu et al., 2022; Wang et al., 2023) which also include personal data (Huang et al., 2022; Kim et al., 2024). But why do some LLMs store personal data? Training LLMs relies on large web-scraped datasets such as Common Crawl (Crawl, 2008) or The Pile (Gao et al., 2020), see also (Villalobos et al., 2024) for a general discussion. Since the internet contains a lot of personal data, this information becomes part of the datasets as well. Memorization of personal data also scales with the capacity of the model (Lu et al., 2024). While on the one hand the memorization is not perfect (Elazar et al., 2021; Cao et al., 2021), it is possible to alter specific facts in LLMs (Meng et al., 2022; Peng et al., 2023).
@@
Legal Perspective. The relationship between AI memorization and data protection law under the GDPR has garnered considerable attention (Wachter et al., 2017; Hacker, 2021; Hacker et al., 2023; Cooper & Grimmelmann, 2024). For this reason, we focus on the GDPR in this paper. While issues of memorization are also highly relevant in the context of copyright law–as highlighted by the New York Times’ lawsuit against Microsoft and OpenAI (Cooper & Grimmelmann, 2024; Sag, 2023; Pope, 2024)–the data protection implications of memorization are the primary focus of this paper. Since LLMs are commonly trained on publicly available internet data, they inevitably pick up personal information. Notably, personal data does not lose its legal status simply because it has been made public. Thus, if this training data is memorized by a model through training, the model itself could potentially be considered personal data under the GDPR (Veale et al., 2018; Juliussen et al., 2023; EDPB, 2024) and trigger unexpected legal consequences. This debate has gained international traction with both the Hamburg Data Protection Commissioner and the Danish Data Protection Authority asserting that LLMs do not contain personal data (Hamburg, 2024; Datatilsynet, 2023), arguing that the data is transformed into abstract mathematical representations and probability weights. However, this position has sparked considerable controversy (Coyer, 2024). The European Data Protection Board [^3] The European Data Protection Board is an independent EU body that ensures the consistent application of the GDPR and provides related guidance, see https://www.edpb.europa.eu/about-edpb/who-we-are/european-data-protection-board_en for more details . has joined the discussion, issuing an opinion stating that AI models trained on personal data cannot automatically be considered anonymous, implying that in most cases, they must be classified as personal data. From our perspective, this stance seems to stem from policy considerations that do not align with the current legal landscape. Instead, it appears to be driven by concerns over the impracticality or undesirability of the legal implications under data protection law. This conflation of technical facts and legal implications has led to misleading conclusions that fail to address the nuanced complexities of ML systems.
@@
ML Research should acknowledge legal implications of LLMs as personal data. The current literature lacks a comprehensive paper that consolidates the various opinions and arguments as to whether LLMs [^4] Please note that in our work, as in the previous (legal) literature, we focus on LLMs. However, our arguments also apply to other modalities, such as Vision Language Models. qualify as personal data under the GDPR. Furthermore, from our point of view, legal implications of this classification remain underexplored, and many developers and computer scientists are unaware of the severe legal implications this issue entails. Therefore, the goal of this paper is twofold: first, we aim to clarify the legal status of LLMs trained on personal data under the GDPR. Second, we argue that ML researchers must acknowledge the legal implications of LLMs as personal data and consider these circumstances during model development. To facilitate this understanding, we contribute the following:
@@ 2.1 Personal Data and Responsibilities of Stakeholders
The GDPR forms the central legal framework for the protection of individuals with regard to the processing of their personal data within the European Union (Jones & Kaminski, 2020). According to Article 4(1) GDPR, personal data is defined as “any information relating to an identified or identifiable natural person.” In other words, if information can be linked back to a person, whether directly or indirectly, it is protected under the GDPR (Finck & Pallas, 2020). Typical examples that first come to mind for personal data in the context of ML applications are names or dates of birth. However, the Court of Justice of the European Union (CJEU) clarified that personal data is “not restricted to information that is sensitive or private, but potentially encompasses all kinds of information, not only objective but also subjective.” [^5] Case C-434/16 Nowak [2017] ECLI:EU:C:2017:994, para 34. This interpretation underscores the broad scope of personal data (Purtova, 2018) and demonstrates that even information that may not initially appear to relate to an individual can still qualify as personal data. For instance, even parameters within a model that encode seemingly anonymized patterns—such as weights linked to frequent phrases in text or common visual features—may still be considered personal data.
@@
The GDPR distinguishes between two key roles: the data subject and the data controller. The data subject is the natural person whose personal data is processed, [^6] Article 4(1) GDPR. e.g., the person whose information is stored in the LLM. The data controller, on the other hand, is the entity that determines the purposes and means of processing the data. [^7] Article 4(7) GDPR. In ML research, determining the data controller is not straightforward. Could it be the principal investigator, the university, or even the ML researcher? Typically, organizations such as companies or institutions act as data controllers because they define how and whether personal data should be processed by their employees. [^8] Case C-131/12 Google Spain and Google [2014] ECLI:EU:C:2014:317, para. 33. They have significant authority over data processing decisions, which qualifies them as data controllers under the GDPR. In the context of ML research, however, the situation is different. Researchers generally enjoy considerable independence due to their constitutionally protected freedom of science and teaching. [^9] See for EU researchers: Article 13 Charter of Fundamental Rights of the European Union. This independence allows them to make their own decisions about the purposes, sources, and methods of processing personal data. Whoever autonomously decides, for example, which datasets to use for training or how to pre-process the data, may qualify as a data controller under the GDPR. Therefore, it has been argued that researchers or principal investigators can be data controllers as well (Research & Innovation, 2024).
@@
While the data subject holds rights, the data controller is bound by obligations. The GDPR grants several rights to data subjects, such as the right to access their data, to request its erasure (the “right to be forgotten”), to object to its processing, and to have their data provided in a portable, machine-readable format (Vrabec, 2021; Wolters, 2018). Conversely, data controllers must ensure data confidentiality, notify authorities of breaches, conduct risk assessments, and process data only when there is a lawful basis, such as consent, and for a specific purpose and limited duration (Hintze, 2018).
@@ 2.2 When Does Data Relate to an Individual?
Under the GDPR, information is considered personal data if it relates to an identified or identifiable person. A person is “identified” when their identity can be directly determined from the information itself. [^10] Case C-582/14 Breyer [2016] ECLI:EU:C:2016:779, para 38. For example, details like a name, date of birth, address, or medical history clearly identify the individual if memorized in an LLM. Moreover, a person is considered “identifiable” if–while the information alone is not sufficient to determine their identity–it becomes possible to do so when combined with additional data (Article 4 (2) GDPR). For example, a dataset that contains browser histories without names might still be personal data. When combined with IP addresses or patterns of visited sites, it can reveal an individual’s identity. Again, if identifiable data are memorized in an LLM, the model itself must be seen as personal data. To determine whether someone is identifiable, the GDPR requires considering all means likely to be used by the data controller or any other party to identify the individual (Recital (26) GDPR), see also Finck & Pallas (2020) for further discussion. This assessment must include factors such as the cost, time, and resources needed for identification, as well as the current state of technology and potential future technological advancements (Finck & Pallas, 2020).
@@ 2.3 The GDPR Trigger: Data Processing
Many rights and obligations of the GDPR are triggered whenever personal data is “processed”. “Processing” is defined as any operation performed on personal data (see Article 4(2) GDPR). Article 4(2) GDPR lists examples of processing activities, including collection, organization, structuring, storage, retrieval, use, disclosure, erasure, or destruction. If personal data can be retrieved from an LLM that memorized it during training within its parameters, the model itself should be classified as personal data. Consequently, any action performed on the model, such as training, uploading, downloading, storing, or deploying it on platforms like GitHub or HuggingFace, fine-tuning, or otherwise sharing it, constitutes processing of personal data.
@@ 3 The Implications of Memory in LLMs
ML developers who act as data controllers under the GDPR (see Section 2.1) must comply with its provisions whenever personal data is used for training. But what happens after training is complete? Does the LLM memorize personal data in a way that it remains “stored” within the model? If so, this raises the critical question of whether a fully trained model should itself be classified as personal data and therefore fall within the scope of the GDPR (EDPB, 2024; Veale et al., 2018; Leiser & Dechesne, 2020; Juliussen et al., 2023). This would trigger significant implications and challenges that can be difficult to comply with in practice (see Section 4). In this section, we briefly discuss both questions and illustrate why LLMs qualify as personal data.

### v2

@@ Abstract
Does GPT know you? The answer depends on your level of public recognition; however, if your information was available on a website, the answer could be yes. Most Large Language Models (LLMs) memorize training data to some extent. Thus, even when an LLM memorizes only a small amount of personal data, it typically falls within the scope of data protection laws. If a person is identified or identifiable, the implications are far-reaching. The LLM is subject to EU General Data Protection Regulation requirements even after the training phase is concluded. To back our arguments: (1.) We reiterate that LLMs output training data at inference time, be it verbatim or in generalized form. (2.) We show that some LLMs can thus be considered personal data on their own. This triggers a cascade of data protection implications such as data subject rights, including rights to access, rectification, or erasure. These rights extend to the information embedded within the AI model. (3.) This paper argues that machine learning researchers must acknowledge the legal implications of LLMs as personal data throughout the full ML development lifecycle, from data collection and curation to model provision on e.g., GitHub or Hugging Face. (4.) We propose different ways for the ML research community to deal with these legal implications. Our paper serves as a starting point for improving the alignment between data protection law and the technical capabilities of LLMs. Our findings underscore the need for more interaction between the legal domain and the ML community.
@@ 1 Introduction
GPT might know you— at least if there is a significant amount of online content about you. Most Large Language Models (LLMs) and corresponding transformer models memorize at least some portions — even if only a very small amount — of the input data provided during training, either in a verbatim or quasi-verbatim manner. [11, 73, 77]. This memorization is double-edged. The training objective of an LLM is partly generalization, but memorization is also an essential component [74, 64, 36, 4]. Without memorization, an LLM would be, for example, unable to tell us that the Eiffel Tower is located in Paris. Thus, memorization is an important and — to some extent — deliberate aspect of training LLMs. While memorization is a crucial feature of LLMs, it also comes with challenges [3], while a key problematic aspect is the consideration of memorized personal data.
@@
Although the problematic nature of personal data memorization in LLMs has been mentioned numerous times, there are debates about the exact amount of memorization [^1] Please note that for our position, the exact amount of memorization is irrelevant. As we discuss later, memorizing even tiny amounts of personal data could trigger the applicability of the GDPR. [9, 39, 11, 89, 14, 42]. The debate about its implications is still in its early stages. While much of the technical work focuses on unlearning techniques [45, 6, 90], the legal implications are often overlooked in the technical ML community [^2] We use the term “ML community” to refer to researchers — whether from academia or industry — who work on topics suitable for publication at conferences like ICML, NeurIPS, or ICLR.. For example, the processing of personal data requires a legal basis. Furthermore, individuals whose data has been memorized by an LLM may have the right to request its erasure from the model, which could require the removal of that information from the model itself, see discussion [79, 86, 27, 1, 45]. These questions are already unfolding in real life. For instance, complaints have been filed against OpenAI, alleging General Data Protection Regulation (GDPR) [^3] EU Regulation 2016/679, 27.4.2016. violations due to the company’s failure to ensure the accuracy of personal data, comply with access and rectification requests, and provide transparency about the sources and processing of personal data in ChatGPT [58]. From our perspective, not all model developers fully consider or acknowledge the serious implications of this requirement.
@@
Why should ML researchers care whether LLMs are personal data or not? If LLMs qualify as personal data, ML researchers may be responsible for their processing. In such cases, researchers must also comply with data protection regulations. As we demonstrate below, non-compliance with data protection regulations can lead to severe fines, whether on a personal, university, or company level. Please also note that some data protection laws, such as the GDPR, even extend to researchers outside the EU as long as they process data from people within the EU (see Article 3 GDPR). Beyond legal compliance, successful LLM research (and its applications) requires not only technical rigor but also the trust of users. We believe that adhering to data protection regulations can enhance trust in ML systems. When researchers commit to data protection safeguards, their applications are more likely to gain public acceptance.
@@
Computational Perspective. LLMs clearly memorize some input data verbatim in normal usage contexts while still generalizing to new contexts, however, there is a debate about the exact amount of memorization [9, 36, 74, 4]. This memorization increases with the growing capacity of the model, the frequency of examples in the training data, and the number of tokens of context used to prompt the model [11]. If data are not outputted verbatim, in some settings, an adversarial attacker is able to extract large amounts of training data [56]. However, it has also been noted that verbatim memorization of facts is an important feature — and not a bug — of LLMs [4]. A weaker form of verbatim memorization is the knowledge of facts in LLMs [43]. LLMs store certain facts [62, 40, 67, 88, 83] which also include personal data [41, 47]. But why do some LLMs store personal data? Training LLMs relies on large web-scraped datasets such as Common Crawl [17] or The Pile [31], see also [78] for a general discussion. Since the internet contains a lot of personal data, this information becomes part of the datasets as well. Memorization of personal data also scales with the capacity of the model [53]. While on the one hand the memorization is not perfect [25, 8], it is possible to alter specific facts in LLMs [54, 60].
@@
Legal Perspective. The relationship between AI memorization and data protection law under the GDPR has garnered considerable attention [82, 33, 34, 14]. For this reason, we focus on the GDPR in this paper. While issues of memorization are also highly relevant in the context of copyright law–as highlighted by the New York Times’ lawsuit against Microsoft and OpenAI [14, 69, 63]–the data protection implications of memorization are the primary focus of this paper. Since LLMs are commonly trained on publicly available internet data, they inevitably pick up personal information. Notably, personal data does not lose its legal status simply because it has been made public [72]. Thus, if this training data is memorized by a model through training, the model itself could potentially be considered personal data under the GDPR [76, 45, 24] and trigger unexpected legal consequences. This debate has gained international traction with both the Hamburg Data Protection Commissioner and the Danish Data Protection Authority asserting that LLMs do not contain personal data [35, 20], arguing that the data is transformed into abstract mathematical representations and probability weights. However, this position has sparked considerable controversy [16]. The European Data Protection Board [^4] The European Data Protection Board is an independent EU body that ensures the consistent application of the GDPR and provides related guidance, see https://www.edpb.europa.eu/about-edpb/who-we-are/european-data-protection-board_en for more details . has joined the discussion, issuing an opinion stating that AI models trained on personal data cannot automatically be considered anonymous, implying that in most cases, they must be classified as personal data. From our perspective, this stance seems to stem from policy considerations that do not align with the current legal landscape. Instead, it appears to be driven by concerns over the impracticality or undesirability of the legal implications under data protection law. This conflation of technical facts and legal implications has led to misleading conclusions that fail to address the nuanced complexities of ML systems.
@@
Our Position: ML Research Must Acknowledge Legal Implications of LLMs as Personal Data. ML Research should acknowledge legal implications of LLMs as personal data. The current literature lacks a comprehensive paper that consolidates the various opinions and arguments as to whether LLMs [^5] Please note that in our work, as in the previous (legal) literature, we focus on LLMs. However, our arguments also apply to other modalities, such as Vision Language Models. qualify as personal data under the GDPR. Furthermore, from our point of view, legal implications of this classification remain underexplored, and many developers and computer scientists are unaware of the severe legal implications this issue entails. Therefore, the goal of this paper is twofold: first, we aim to clarify the legal status of LLMs trained on personal data under the GDPR. Second, we argue that ML researchers must acknowledge the legal implications of LLMs as personal data and consider these circumstances during model development. To facilitate this understanding, we contribute the following:
@@ 2.1 Personal Data and Responsibilities of Stakeholders
The GDPR forms the central legal framework for the protection of individuals with regard to the processing of their personal data within the European Union [44]. According to Article 4(1) GDPR, personal data is defined as “any information relating to an identified or identifiable natural person.” In other words, if information can be linked back to a person, whether directly or indirectly, it is protected under the GDPR [30]. Typical examples that first come to mind for personal data in the context of ML applications are names or dates of birth. However, the Court of Justice of the European Union (CJEU) clarified that personal data is “not restricted to information that is sensitive or private, but potentially encompasses all kinds of information, not only objective but also subjective.” [^6] Case C-434/16 Nowak [2017] ECLI:EU:C:2017:994, para 34. This interpretation underscores the broad scope of personal data [65] and demonstrates that even information that may not initially appear to relate to an individual can still qualify as personal data. For instance, even parameters within a model that encode seemingly anonymized patterns—such as weights linked to frequent phrases in text or common visual features—may still be considered personal data.
@@
The GDPR distinguishes between two key roles: the data subject and the data controller. The data subject is the natural person whose personal data is processed, [^7] Article 4(1) GDPR. e.g., the person whose information is stored in the LLM. The data controller, on the other hand, is the entity that determines the purposes and means of processing the data. [^8] Article 4(7) GDPR. In ML research, determining the data controller is not straightforward. Could it be the principal investigator, the university, or even the ML researcher? Typically, organizations such as companies or institutions act as data controllers because they define how and whether personal data should be processed by their employees. [^9] Case C-131/12 Google Spain and Google [2014] ECLI:EU:C:2014:317, para. 33. They have significant authority over data processing decisions, which qualifies them as data controllers under the GDPR. In the context of ML research, however, the situation is different. Researchers generally enjoy considerable independence due to their constitutionally protected freedom of science and teaching [^10] See for EU researchers: Article 13 Charter of Fundamental Rights of the European Union..This independence allows them to make their own decisions about the purposes, sources, and methods of processing personal data. Whoever autonomously decides, for example, which datasets to use for training or how to pre-process the data, may qualify as a data controller under the GDPR. Therefore, it has been argued that researchers or principal investigators can be data controllers as well [66].
@@
While the data subject holds rights, the data controller is bound by obligations. The GDPR grants several rights to data subjects, such as the right to access their data, to request its erasure (the “right to be forgotten”), to object to its processing, and to have their data provided in a portable, machine-readable format [80, 84]. Conversely, data controllers must ensure data confidentiality, notify authorities of breaches, conduct risk assessments, and process data only when there is a lawful basis, such as consent, and for a specific purpose and limited duration [37].
@@ 2.2 When Does Data Relate to an Individual?
Under the GDPR, information is considered personal data if it relates to an identified or identifiable person. A person is “identified” when their identity can be directly determined from the information itself. [^11] Case C-582/14 Breyer [2016] ECLI:EU:C:2016:779, para 38. For example, details like a name, date of birth, address, or medical history clearly identify the individual if memorized in an LLM. Moreover, a person is considered “identifiable” if–while the information alone is not sufficient to determine their identity–it becomes possible to do so when combined with additional data (Article 4 (2) GDPR). For example, a dataset that contains browser histories without names might still be personal data. When combined with IP addresses or patterns of visited sites, it can reveal an individual’s identity. Again, if identifiable data are memorized in an LLM, the model itself must be seen as personal data. To determine whether someone is identifiable, the GDPR requires considering all means likely to be used by the data controller or any other party to identify the individual (Recital (26) GDPR), see also Finck & Pallas [30] for further discussion. This assessment must include factors such as the cost, time, and resources needed for identification, as well as the current state of technology and potential future technological advancements [30].
Counterintuitively, data from public figures are also considered personal data and thus trigger the applicability of the GDPR [72]. For the applicability of the GDPR, it does not matter whether personal data is private or public. However, it could affect which legal basis can be used for processing the personal data (see section 4).
@@ 2.3 The GDPR Trigger: Data Processing
Many rights and obligations of the GDPR are triggered whenever personal data is “processed”. “Processing” is defined as any operation performed on personal data (see Article 4(2) GDPR). Article 4(2) GDPR lists examples of processing activities, including collection, organization, structuring, storage, retrieval, use, disclosure, erasure, or destruction. If personal data can be retrieved from an LLM that was memorized during training within its parameters, the model itself should be classified as personal data. Consequently, any action performed on the model, such as training, uploading, downloading, storing, or deploying it on platforms like GitHub or HuggingFace, fine-tuning, or otherwise sharing it, constitutes processing of personal data.
Under the GDPR, responsibility for data processing does not lie solely with the entity closest to the data collection. Rather, every single data processing activity, such as storing, accessing, or deleting personal data, requires a legal basis. Accordingly, each entity that processes personal data is independently responsible for ensuring that its specific processing activity complies with the GDPR. Unless special cases apply, such as order processing (Article 28 GDPR) or joint responsibility (Article 26 GDPR), each entity may be held independently liable.
@@ 3 The Implications of Memory in LLMs
ML developers who act as data controllers under the GDPR (see Section 2.1) must comply with its provisions whenever personal data is used for training. But what happens after training is complete? Does the LLM memorize personal data in a way that it remains “stored” within the model? If so, this raises the critical question of whether a fully trained model should itself be classified as personal data and therefore fall within the scope of the GDPR [24, 76, 51, 45]. This would trigger significant implications and challenges that can be difficult to comply with in practice (see Section 4). In this section, we briefly discuss both questions and illustrate why LLMs qualify as personal data.